# Mimicking Human Intuition: Cognitive Belief-Driven Reinforcement Learning

### v3

@@ B.1 Lemma
Lemma B.2 (Convergence of Smoothed Bellman Operator).
@@
B.2 Smoothing Bellman Strategy
@@
         Softmax          $b_{t}=\frac{e^{Q(s,a)}}{\sum_{b}e^{Q(s,b)}}$
         Clipped Max           $b_{t}=\begin{cases}1-\tau,&\text{if }a=a^{*}\\ \frac{\tau}{A-1},&\text{if }a\neq a^{*}\end{cases}$
         Clipped Softmax           $b_{t}=\begin{cases}\frac{e^{\beta Q(s,a)}}{\sum_{b\in I}e^{\beta Q(s,b)}},&\text{if }a\in I\\ 0,&\text{if }a\notin I\end{cases}$
         Bayesian Inference           $Q_{\text{adjusted}}(s,a)=Q(s,a)+\mu_{\text{prior}}$ $b_{t}=\frac{e^{Q_{\text{adjusted}}(s^{\prime},a)}}{\sum_{b}e^{Q_{\text{adjusted}}(s^{\prime},b)}}$ $\sigma_{\text{posterior}}^{2}=\left(\frac{1}{\sigma_{\text{prior}}^{2}}+\frac{n}{\sigma_{\text{observation}}^{2}}\right)^{-1}$ $\mu_{\text{posterior}}=\sigma_{\text{posterior}}^{2}\left(\frac{\mu_{\text{prior}}}{\sigma_{\text{prior}}^{2}}+\sum_{i=1}^{n}\frac{r_{i}}{\sigma_{\text{observation}}^{2}}\right)$
@@
Appendix C Convergence Proof
We outline a proof that builds upon the following result (Singh et al., 2000; Barber, 2023) and follows the framework provided in (Melo, 2001):
Theorem C.1.
@@ Appendix C Convergence Proof
$=(1-\alpha_{t})\Delta_{t}+\alpha_{t}\left(r_{t}+\gamma\langle Q(s_{t+1},a)\rangle_{a}-Q_{*}(s_{t},a_{t})\right)$ (28)
@@
Since the rewards $r_{t}$ are bounded ($|r_{t}|\leq R_{\text{max}}$) and the discount factor $10<\gamma<$ 1, we can show that $Q_{t}$ remains bounded independently of the convergence of $\Delta_{t}$.
@@
Since $10<\gamma<$ 1, the infinite sum $\sum_{k=0}^{\infty}\gamma^{k}$ is a geometric series that sums to:
@@
Appendix D Experiment Setting
D.1 Classic Control and Box 2D Environment
@@
D.2 MetaDrive Block Type Description
@@ X Intersection
D.3 Map Design and Testing Objectives
D.3.1 Map 1: SrOYCTRyS
@@ D.5 Atari Environments
D.6 Environment Parameter & Agent Parameter
Table 4: Q-family vs PPO Algorithm and Environment Parameters
Parameter Q-Family PPO
Discrete Action Space True
Policy Basic_Q_network Categorical_AC
Representation Basic_MLP
Runner DRL
Representation Hidden Size [256, 256] [512,]
Q/Actor Hidden Size [256, 256] [256, 256]
Critic Hidden Size N/A [256, 256]
Activation Function relu leaky_relu
Activation for Actions N/A tanh
Seed 123 / 321 / 666
Number of Parallels 10
Buffer Size 500,000 Horizon_Size * Parallels (128 * 10)
Batch Size 64 N/A
Horizon Size N/A 128
Number of Epochs N/A 4
Number of Minibatches N/A 4
Learning Rate 0.00025
Start Greedy 1.0 N/A
End Greedy 0.01 N/A
Decay Step for Greedy 50,000 N/A
Sync Frequency 50 N/A
Training Frequency 1 N/A
Start Training Step 1,000 N/A
Running Steps 2,000,000
Use Gradient Clipping N/A True
Value Function Coefficient N/A 0.25
Entropy Coefficient N/A 0.0
Target KL Divergence N/A 0.001
Clip Range N/A 0.2
Clip Gradient Norm N/A 0.5
Gamma 0.99
Use GAE N/A True
GAE Lambda N/A 0.95
Use Advantage Normalization N/A True
Use Observation Normalization False True
Use Reward Normalization False True
Observation Normalization Range 5
Reward Normalization Range 5
Test Steps 10,000
Evaluation Interval 50,000 5,000
Test Episodes 5
Appendix E Running Setting
For the Cartpole and Lunar Lander environments, the training process utilizes 1 RTX 3060 and typically runs less than 30 minutes. For the Carracing environment, we require 1 RTX 3060 and 2 hours of running. For the Metadrive environments, the training process utilizes 1 RTX 3060 and typically runs around 3-6 hours according to different complexity.

### v4

@@ A.3 Smoothed Bellman Operator
Lemma 28 establishes that the Smoothed Bellman Operator provides a conservative backup: replacing the hard maximization with a belief-weighted expectation yields an update that forms a lower bound on the classical Bellman backup, thereby stabilizing value propagation under uncertainty.
Lemma A.2 (Convergence of Smoothed Bellman Operator).
@@
Lemma A.2 complements this by showing that if the action-preference distribution $q_{t}(\cdot\mid s_{t+1})$ asymptotically collapses onto the optimal action, then iterative application of the Smoothed Bellman Operator converges to the optimal value function $Q^{*}$. Together, these results establish that the Smoothed Bellman Operator not only smooths value backups for improved robustness, but also preserves the fundamental convergence guarantees of classical reinforcement learning. The full proof of Lemma A.2 is provided in next subsection.
To instantiate the Smoothed Bellman Operator in practice, different smoothing strategies can be employed to construct the action-preference distribution $b_{t}$. These strategies determine how strongly the update deviates from hard maximization and how uncertainty is incorporated. Representative examples are summarized in Table 3.
@@
    Softmax     $q_{t}=\frac{e^{Q(s,a)}}{\sum_{b}e^{Q(s,b)}}$
    Clipped Max      $q_{t}=\begin{cases}1-\tau,&\text{if }a=a^{*}\\ \frac{\tau}{A-1},&\text{if }a\neq a^{*}\end{cases}$
    Clipped Softmax      $q_{t}=\begin{cases}\frac{e^{\beta Q(s,a)}}{\sum_{b\in I}e^{\beta Q(s,b)}},&\text{if }a\in I\\ 0,&\text{if }a\notin I\end{cases}$
@@
A.3.1 Convergence Proof
We outline a proof that builds upon the following result (singh2000convergence; barber2023smoothed) and follows the framework provided in (melo2001convergence):
Theorem A.3.
@@ A.3.1 Convergence Proof
$(1-\alpha_{t})\Delta_{t}+\alpha_{t}\left(r_{t}+\gamma\langle Q(s_{t+1},a)\rangle_{a}-Q_{*}(s_{t},a_{t})\right)$ (34)
@@
Since the rewards $r_{t}$ are bounded ($|r_{t}|\leq R_{\text{max}}$) and the discount factor $0<\gamma<1$, we can show that $Q_{t}$ remains bounded independently of the convergence of $\Delta_{t}$.
@@
Since $0<\gamma<1$, the infinite sum $\sum_{k=0}^{\infty}\gamma^{k}$ is a geometric series that sums to:
@@
A.4 Experiment Setting
A.4.1 Classic Control and Box 2D Environment
@@
A.4.2 MetaDrive Block Type Description
@@ X Intersection
A.4.3 Map Design and Testing Objectives
Map 1: SrOYCTRyS